# Byzantine Consensus in the Random Asynchronous Model

### v1

@@ Abstract.
1. Introduction
Byzantine Fault-Tolerant (BFT) consensus is a fundamental primitive in distributed computing, serving as the backbone of many applications, including blockchains (Bano et al., 2019). Solving BFT consensus in an asynchronous network model (Fischer et al., 1983) is a well-studied problem (Cachin et al., 2011; Keidar et al., 2021; Ren et al., 2017; Jovanovic et al., 2024). Typically, algorithms in this model assume that the adversary controls (1) a subset of faulty processes and (2) message delays, particularly the message schedule—the order in which messages are delivered.
However, asynchronous BFT consensus is constrained by restrictive lower bounds (Fischer et al., 1983; Lamport et al., 1982). These lower bounds often stem from the adversary’s ability to impose arbitrary message schedules. Requiring an algorithm to function under the worst possible scheduling scenario often renders tasks impossible due to a single, highly unlikely counterexample, that is, a pathological schedule that violates the algorithm’s properties (Fischer et al., 1983).
Yet, due to performance considerations, practical implementations of consensus circumvent the limitations of asynchrony by assuming that the schedule is not adversarial. For example, we are aware of asynchronous consensus protocols running in production commodity wide-area networks for extensive periods of time without a random coin implementation (Danezis et al., 2022)—so as to avoid the performance overhead of generating cryptographically-secure randomness—without loss of liveness. This motivates our search for a model that explains this empirical observation.
Another common relaxation that circumvents the limitations of asynchrony is to assume periods of synchrony, leading to the widely used partially synchronous model (Dwork et al., 1988). However, protocols based on partial synchrony lose liveness outside of these periods, which may occur due to poor network conditions or denial-of-service (DoS) attacks (Giuliari et al., 2024).
@@ 1. Introduction
To explore this question, we introduce a new variant of the asynchronous model, which we call the random asynchronous model. In this model, message delays remain unbounded, and the adversary still controls a subset of processes. However, the message schedule is no longer adversarial; instead, messages are delivered in a random order. By relaxing adversarial scheduling, our approach unlocks new algorithmic possibilities for consensus. It enables protocols to achieve probabilistic guarantees (Avelas et al., 2024) that were previously impossible in the standard asynchronous model, while preserving unbounded message delays and Byzantine faults.
@@
Our model enables consensus protocols that are impossible under standard asynchrony. The key insight is that the random asynchronous model prevents an adversary from blocking honest parties from communicating indefinitely. In traditional asynchrony, an adversary can delay messages indefinitely to prevent termination (e.g., in the FLP impossibility result (Fischer et al., 1983)). In contrast, in the random asynchronous model, the adversary cannot control the schedule, ensuring that honest parties can exchange messages within a bounded number of steps with high probability.
1.1. Modeling Challenges
@@
A natural concern is whether removing adversarial scheduling trivializes the consensus problem. We address this in two ways. First, Appendix A.2 presents a naive algorithm that fails to solve consensus when $n\leq 2f+1$, highlighting a fundamental challenge: even though the random asynchronous model guarantees eventual communication, Byzantine nodes can still equivocate, preventing honest processes from distinguishing between correct and faulty messages. Second, we complement our positive results with corresponding impossibility results, establishing close bounds on what can be achieved in the random asynchronous model.
1.2. Our Results
@@
Resilience Positive Result Negative Result
$n=3f+1$ Det. strong V, det. A, P1 T Det. strong V, det. A, det. T
$n=2f+1$ Whp strong V, whp A, det. T P1 strong V, P1 A, det. T
$n=f+2$ Det. weak V, whp A, det. T Whp strong V, whp A, det. T
Table 1. Results in this paper for Byzantine consensus. Det. means “deterministic”, P1 means “with probability 1”, and whp means “with high probability”. V stands for “validity”, A stands for “agreement”, and T stands for “termination”. These terms are defined in Section 2. Additional assumption needed (see Section 6).
2. Model & Prelimnaries
Processes
@@ 2. Model & Prelimnaries
We consider a set $\Pi$ of $n$ processes, up to $f$ of which may be faulty. We consider the Byzantine-fault model, in which faulty processes may depart arbitrarily from the protocol. Throughout the paper, we assume that correct (non-faulty) processes have deterministic logic, i.e., they do not have access to a source of randomness.
Cryptography
@@
We make standard assumptions: processes communicate through authenticated channels and have access to digital signatures whose properties cannot be broken by the adversary.
Network
@@
The processes communicate by sending messages over a fully-connected reliable network: every pair of processes $p$ and $q$ communicate over a link that satisfies the integrity and no-loss properties. Integrity requires that a message $m$ from $p$ be received by $q$ at most once and only if $m$ was previously sent by $p$ to $q$. No-loss requires that a message $m$ sent from $p$ to $q$ be eventually received by $q$.
Random asynchrony
@@
Processes send messages by submitting them to the network; a random scheduler decides in which order submitted messages are delivered. We assume the scheduler delivers messages one by one, i.e., no two delivery events occur at exactly the same time. For convenience, we use a global discrete notion of time to reflect the sequence of delivery events: time proceeds in discrete steps $0,1,\ldots$; each time step corresponds to a message delivery event in the system. Processes do not have access to this notion of time (they do not have clocks). Sending a message and performing local computation occur instantaneously, between time steps. Note that this notion of time does not bound message delays: an arbitrary amount of real time can pass between time steps.
@@
We next describe the random scheduler. At any time $t$, let $P(t)\subseteq\Pi^{2}$ be the set of pairs of processes $(p,q)$ such that $p$ has at least one pending message to $q$. We say that a message $m$ from $p$ to $q$ is pending at time $t$ if $p$ send $m$ before $t$ and $q$ has not received $m$ by time $t$. At each time step $t$, the random scheduler draws a pair $(p,q)$ at random from $P(t)$ and delivers to $q$ the earliest message from $p$. We assume that there exists a constant $\mathcal{C}(n,f)>0$ such that, for any $p$ and $q$, the probability that $(p,q)\in P(t)$ is drawn at time $t$ is at least $\mathcal{C}(n,f)$. In general, $\mathcal{C}(n,f)$ may depend on $n$ and $f$, but not on $t$. We assume that each scheduling draw is independent from others, and does not depend on the content of the message being delivered.
Consensus
@@
Our algorithms solve two variants of binary Byzantine consensus: strong and weak. Strong Byzantine consensus is defined by the following properties:
Strong Validity:
@@
Weak Byzantine consensus (Lamport, 1983) has the same agreement and termination properties as the strong variant above, but has a different validity property:
Weak Validity:
@@
Deterministic and probabilistic properties
The probability of a schedule is the probability of the intersection of all its scheduling steps. We say that an algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ deterministically if $\mathcal{P}$ holds in every execution of $\mathcal{A}$. Note that any algorithm that ensures a property $\mathcal{P}$ deterministically in the standard asynchronous model, must also ensure $\mathcal{P}$ deterministically in the random asynchronous model. Given an algorithm $\mathcal{A}$ and a property $\mathcal{P}$, we say that a schedule $S$ is bad for $\mathcal{P}$ if there exists an execution $E$ of $\mathcal{A}$ with schedule $S$ such that $\mathcal{P}$ does not hold in $E$. An algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ with probability $1$ if the total probability of all schedules that are bad for $\mathcal{P}$ is 00. An algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ with high probability (whp) if the total probability of all schedules that are bad for $\mathcal{P}$ is negligible; in this paper, a probability is negligible if approaches 00 exponentially with some parameter of the algorithm, for any (fixed) $n$ and $f$ (e.g., the number of communication steps).
3. $n=3f+1$: Deterministic safety, termination with probability 1
In this section we solve binary Byzantine consensus with deterministic strong validity and agreement, and termination with probability $1$.
Our algorithm, shown in Algorithm 1, proceeds in rounds: in each round, processes attempt to decide using a Round procedure; if unsuccessful, processes update their estimate and try again in the next round. The Round procedure is similar to an adopt-commit object (Gafni, 1998; Yang et al., 1998): processes propose a value and return a pair $(g,v)$, where $v$ is a value and $g$ is a grade, which can be either Commit or Adopt. If $g=\textsc{Commit}$, then it is guaranteed that all correct processes return the same value $v$ (possibly with different grades). If all correct processes propose the same value $v$, then all correct processes are guaranteed to return $v$ with a Commit grade.
@@ 3. $n=3f+1$: Deterministic safety, termination with probability 1
6:   if $\exists\,v^{*}$ such that I received $>n/2$ $\langle\textsc{Init},r,v^{*}\rangle$ messages:
@@
7:   BRB-Broadcast $\langle\textsc{Echo},r,proposal,\mathcal{H}\rangle$ to all processes $\triangleright$ Phase 2
@@
8:   Wait to BRB-Deliver valid $\langle\textsc{Echo},r,\_,\_\rangle$ messages from $n-f$ processes
@@
9:   if $\exists\,v^{*}\neq\bot$ such that I received $\geq 2f+1$ $\langle\textsc{Echo},r,v^{*},\_\rangle$ messages:
10:      return $\langle\textsc{Commit},v^{*}\rangle$
11:   else:
@@
12:      $v^{*}\leftarrow$ majority value among received $\langle\textsc{Echo},r,\_,\_\rangle$ messages
13:      return $\langle\textsc{Adopt},v^{*}\rangle$
@@
The core of the algorithm is the Round procedure, shown in Algorithm 2. There are two types of messages: Init and Echo. Processes rely on Byzantine Reliable Broadcast (BRB) (Cachin et al., 2011) for communication. [^1] Any asynchronous BRB protocol is also correct in the random asynchronous model. Furthermore, all messages are signed. The algorithm has two phases. In phase 1, each correct process $p$ broadcasts (using BRB) its input value $v$ in an Init message (line 2), and waits to deliver $n-f=2f+1$ Init messages (line 3). Process $p$ adopts as its proposal for phase 2, the majority value among the received Init messages (line 5). Then, in phase 2, $p$ broadcasts an Echo message with $p$’s phase 2 proposal, as well as the $n-f$ (signed) Init messages that justify $v$ to be the majority phase 1 value (line 7); $p$ waits for $n-f$ valid Echo messages (line 8). If all delivered Echo messages are for the same value $v^{*}$, then $p$ commits $v^{*}$ (line 10); otherwise $p$ adopts the majority value (line 13).
@@
We prove the correctness of our algorithm in Appendix B. The main intuition is that at each round, a favorable schedule can help all correct processes agree (i.e., adopt the same estimate), by causing them to select the same majority value in phase 1 (e.g., by ensuring that they deliver Init messages from the same set of processes). Since the schedule is random, it has a non-zero chance of being favorable at each round. Thus, with probability $1$, the schedule will eventually be favorable. This is similar in spirit to random coin based Byzantine Randomized Consensus algorithms (Cachin et al., 2011), in which correct processes choose their next estimate by tossing a coin, if they do not manage to reach agreement in a given round. Here we are relying on the random schedule instead of the random coin. However, as we show later, in our model, consensus is solvable for settings where it is impossible for standard asynchrony even when equipped with a common coin.
Crash-fault tolerant consensus
@@
A similar approach can be used to solve crash-fault tolerant consensus with $n=2f+1$, with the same guarantees of deterministic safety, as well as termination with probability $1$. Our algorithm uses the same round-based structure in Algorithm 1 and a modified Round procedure, that does not require reliable broadcast, and employs standard point-to-point messages instead. The main intuition is similar to our Byzantine algorithm: the random scheduler eliminates the need for a common coin by ensuring that correct processes eventually deliver messages in a favorable order which leads them to agree and thus terminate. Appendix C gives our algorithm and proves its correctness.
4. $n=2f+1$: Deterministic termination, safety whp

### v2

@@ Abstract
keywords:
network model, asynchronous, random scheduler, Byzantine consensus
1 Introduction
Byzantine Fault-Tolerant (BFT) consensus is a fundamental primitive in distributed computing, serving as the backbone of many applications, including blockchains [6]. Solving BFT consensus in an asynchronous network model [13] is a well-studied problem [9, 17, 24, 16]. Typically, algorithms in this model assume that the adversary controls (1) a subset of faulty processes and (2) message delays, particularly the message schedule—the order in which messages are delivered.
However, asynchronous BFT consensus is constrained by restrictive lower bounds [13, 19]. These lower bounds often stem from the adversary’s ability to impose arbitrary message schedules. Requiring an algorithm to function under the worst possible scheduling scenario often renders tasks impossible due to a single, highly unlikely counterexample, that is, a pathological schedule that violates the algorithm’s properties [13].
Yet, due to performance considerations, practical implementations of consensus circumvent the limitations of asynchrony by assuming that the schedule is not adversarial. For example, we are aware of asynchronous consensus protocols running in production over commodity wide-area networks for extensive periods of time without a random coin implementation [11]—so as to avoid the performance overhead of cryptographically-secure randomness—without loss of liveness. This motivates our search for a model to explain this empirical observation.
Another common relaxation that circumvents the limitations of asynchrony is to assume periods of synchrony, leading to the widely used partially synchronous model [12]. However, protocols based on partial synchrony lose liveness outside of these periods, which may occur due to poor network conditions or denial-of-service (DoS) attacks [15].
@@ 1 Introduction
To explore this question, we introduce a new variant of the asynchronous model, which we call the random asynchronous model. In this model, message delays remain unbounded, and the adversary still controls a subset of processes. However, the message schedule is no longer adversarial; instead, messages are delivered in a random order. Our aim is not to argue that this model is a more realistic description of real networks; rather, we view it as a useful abstraction, inspired by assumptions made in practice [11], that lets us analyze the non-adversarial schedule assumption and explore which tasks it enables under unbounded delays while preserving Byzantine faults. This work offers a new perspective on the role of adversarial scheduling in the standard asynchronous setting: by removing the adversary’s scheduling power, we explore what becomes achievable and what remains fundamentally impossible. Our work is a part of the broader line of research that explores relaxations of the fully adversarial scheduler: from the fair-scheduler abstraction of Bracha and Toueg [8], through Aspnes’s noisy scheduling for shared memory [4], to recent DAG-based BFT systems such as Tusk [11] and Mahi-Mahi [16].
@@
We find that our model enables consensus protocols that are impossible under standard asynchrony. The key insight is that the random asynchronous model prevents an adversary from blocking honest parties from communicating indefinitely. In traditional asynchrony, an adversary can delay messages indefinitely to prevent termination (e.g., in the FLP impossibility result [13]). In contrast, in the random asynchronous model, the adversary cannot control the schedule, ensuring that honest parties can exchange messages within a bounded number of steps with high probability.
This last property is reminiscent of what synchrony guarantees, i.e., that any message sent by a correct process is received within a time bound $\Delta$. This leads to the natural question: How powerful is the random asynchronous model with respect to the standard models—asynchrony, partial synchrony, and synchrony? We answer this question in Section 7. In brief, we show that the random asynchronous model sits between asynchrony and synchrony in terms of task solvability power, yet perhaps surprisingly, is incomparable to partial synchrony. To illustrate the latter separation, consider the following two tasks: (1) deterministic consensus when at least one process may fail by crashing, and (2) Byzantine consensus for $n<3f+1$. Task (1) is not solvable in the random asynchronous model, as we show in Section 6, but is solvable in partial synchrony [9]; conversely, task (2) is not solvable in partial synchrony [12] yet becomes attainable in the random asynchronous model with probabilistic safety (Section 4). Finally, we also show that the random asynchronous model behaves like a variant of synchrony where timing guarantees hold with high probability instead of deterministically, thus confirming the intuitive similarity between the two models.
1.1 Modeling Challenges
@@
A natural concern is whether removing adversarial scheduling trivializes the consensus problem. We address this in two ways. First, Appendix A.2 presents a naive algorithm that fails to solve consensus when $n\leq 2f+1$, highlighting a fundamental challenge: even though the random asynchronous model guarantees eventual communication, Byzantine nodes can still equivocate, preventing honest processes from distinguishing between correct and faulty messages. Second, we complement our positive results with corresponding impossibility results, establishing close bounds on what is solvable in the random asynchronous model.
1.2 Our Results
@@
Table 1: Guarantees achievable (top) and impossible (bottom) in the random asynchronous model. “Strong” and “weak” validity are defined in Section 2.
$n$ versus $f$ Validity Agreement Termination
Positive results
$n=3f+1$ Deterministic strong Deterministic Almost surely
$n=2f+1$ With high probability strong With high probability Deterministic
$n=f+2$ Deterministic weak With high probability Deterministic
Negative results
$n=3f+1$ Deterministic strong Deterministic Deterministic
$n=2f+1$ Almost surely strong Almost surely Deterministic
$n=f+2$ With high probability strong With high probability Deterministic
2 Model & Preliminaries
@@
Processes. We consider a set $\Pi$ of $n$ processes, up to $f$ of which may be faulty. We consider the Byzantine-fault model, in which faulty processes may depart arbitrarily from the protocol. Throughout the paper, we assume that correct (non-faulty) processes have deterministic logic, i.e., they do not have access to a source of randomness.
@@
Cryptography. We make assumptions that are standard for Byzantine fault-tolerant algorithms: processes communicate through authenticated channels (e.g., using message authentication codes (MACs)) and have access to digital signatures whose properties cannot be broken by the adversary.
@@
Network. The processes communicate by sending messages over a fully-connected reliable network: every pair of processes $p$ and $q$ communicate over a link that satisfies the integrity and no-loss properties. Integrity requires that a message $m$ from $p$ be received by $q$ at most once and only if $m$ was previously sent by $p$ to $q$. No-loss requires that a message $m$ sent from $p$ to $q$ be eventually received by $q$.
@@
Random asynchrony. Processes send messages by submitting them to the network; a random scheduler decides in which order submitted messages are delivered. We assume the scheduler delivers messages one by one, i.e., no two delivery events occur at exactly the same time. For convenience, we use a global discrete notion of time to reflect the sequence of delivery events: time proceeds in discrete steps $0,1,\ldots$; each time step corresponds to a message delivery event in the system. Processes do not have access to this notion of time (they do not have clocks). Sending a message and performing local computation occur instantaneously, between time steps. Note that this notion of time does not bound message delays: an arbitrary amount of real time can pass between time steps.
@@
We next describe the random scheduler. At any time $t$, let $P(t)\subseteq\Pi^{2}$ be the set of pairs of processes $(p,q)$ such that $p$ has at least one pending message to $q$. We say that a message $m$ from $p$ to $q$ is pending at time $t$ if $p$ sends $m$ before $t$ and $q$ has not received $m$ by time $t$. At each time step $t$, the random scheduler draws a pair $(p,q)$ at random from $P(t)$ and delivers to $q$ the earliest message from $p$. We assume that there exists a constant $\mathcal{C}(n,f)>0$ such that, for any $p$ and $q$, the probability that $(p,q)\in P(t)$ is drawn at time $t$ is at least $\mathcal{C}(n,f)$. In general, $\mathcal{C}(n,f)$ may depend on $n$ and $f$, but not on $t$. We assume that scheduling draws do not depend on the content of the message being delivered.
Furthermore, we assume that scheduling draws are conditionally independent of each other, given the history of messages sent and delivered. More formally, let $\{X_{t}\}_{t\geq 0}\in P(t)$ be the random variables that represent the scheduling draws at each step $t$, and let $\mathcal{F}_{t}$ be the history (all messages sent/delivered) up to step $t$. Then our lower bound assumption above can be written as
$\Pr[X_{t}=(p,q)\mid\mathcal{F}_{t}]\;\geq\;C(n,f)\quad\text{for any}\quad t\geq 0$ (1)
and the conditional independence assumption can be written as
$\Pr[X_{t_{1}},\dots,X_{t_{k}}\mid\mathcal{F}_{t_{k}}]=\prod_{i=1}^{k}\Pr[X_{t_{i}}\mid\mathcal{F}_{t_{i}}]\quad\text{for any }t_{1}<\dots<t_{k}.$ (2)
Intuitively, the scheduling draws are independent given the message history because the scheduler flips a fresh coin at every step.
@@
Consensus. Our algorithms solve two variants of binary Byzantine consensus: strong and weak. Strong Byzantine consensus is defined by the following properties:
Strong Validity
@@
Weak Byzantine consensus [18] has the same agreement and termination properties as the strong variant above, but has a different validity property:
Weak Validity
@@
Deterministic and probabilistic properties. The probability of a schedule is the probability of the intersection of all its scheduling steps. We say that an algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ deterministically if $\mathcal{P}$ holds in every execution of $\mathcal{A}$. Note that any algorithm that ensures a property $\mathcal{P}$ deterministically in the standard asynchronous model, must also ensure $\mathcal{P}$ deterministically in the random asynchronous model. Given an algorithm $\mathcal{A}$ and a property $\mathcal{P}$, we say that a schedule $S$ is bad for $\mathcal{P}$ if there exists an execution $E$ of $\mathcal{A}$ with schedule $S$ such that $\mathcal{P}$ does not hold in $E$. An algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ almost surely (a.s.) if the total probability of all schedules that are bad for $\mathcal{P}$ is 00. An algorithm $\mathcal{A}$ ensures a property $\mathcal{P}$ with high probability (w.h.p.) if the total probability of all schedules that are bad for $\mathcal{P}$ is negligible; in this paper, a probability is negligible if it approaches 00 exponentially with some parameter of the algorithm, for any (fixed) $n$ and $f$ (e.g., the number of communication steps).
3 $n=3f+1$: Deterministic Safety, Termination Almost Surely
In this section we solve binary Byzantine consensus with deterministic strong validity and agreement, and termination almost surely.
Our algorithm, shown in Algorithm 1, proceeds in rounds: in each round, processes attempt to decide using a Round procedure; if unsuccessful, processes update their estimate and try again in the next round. The Round procedure is similar to an adopt-commit object [14, 26]: processes propose a value and return a pair $(g,v)$, where $v$ is a value and $g$ is a grade, which can be either Commit or Adopt. If $g=\textsc{Commit}$, then it is guaranteed that all correct processes return the same value $v$ (possibly with different grades). If all correct processes propose the same value $v$, then all correct processes are guaranteed to return $v$ with a Commit grade.
@@ 3 $n=3f+1$: Deterministic Safety, Termination Almost Surely
6:   BRB-Broadcast $\langle\textsc{Echo},r,proposal,\mathcal{H}\rangle$ to all processes $\triangleright$ Phase 2
@@
7:   Wait to BRB-Deliver valid $\langle\textsc{Echo},r,\_,\_\rangle$ messages from $n-f$ processes
@@
8:   if $\exists\,v^{*}\neq\bot$ such that I received $\geq 2f+1$ $\langle\textsc{Echo},r,v^{*},\_\rangle$ messages:
9:      return $\langle\textsc{Commit},v^{*}\rangle$
10:   else:
@@
11:      $v^{*}\leftarrow$ majority value among received $\langle\textsc{Echo},r,\_,\_\rangle$ messages
12:      return $\langle\textsc{Adopt},v^{*}\rangle$
@@
The core of the algorithm is the Round procedure, shown in Algorithm 2. There are two types of messages: Init and Echo. Processes rely on Byzantine Reliable Broadcast (BRB) [9] for communication. [^1] Any asynchronous BRB protocol is also correct in the random asynchronous model. Furthermore, all messages are signed. The algorithm has two phases. In phase 1, each correct process $p$ broadcasts (using BRB) its input value $v$ in an Init message (line 2), and waits to deliver $n-f=2f+1$ Init messages (line 3). Process $p$ adopts as its proposal for phase 2, the majority value among the received Init messages (line 5). Then, in phase 2, $p$ broadcasts an Echo message with $p$’s phase 2 proposal, as well as the $n-f$ (signed) Init messages that justify $v$ to be the majority phase 1 value (line 6); $p$ waits for $n-f$ valid Echo messages (line 7). If all delivered Echo messages are for the same value $v^{*}$, then $p$ commits $v^{*}$ (line 9); otherwise $p$ adopts the majority value (line 12).
@@
We prove the correctness of our algorithm in Appendix B. The main intuition is that at each round, a favorable schedule can help all correct processes agree (i.e., adopt the same estimate), by causing them to select the same majority value in phase 1 (e.g., by ensuring that they deliver Init messages from the same set of processes). Since the schedule is random, it has a non-zero chance of being favorable at each round. Thus, almost surely, the schedule will eventually be favorable. This is similar in spirit to random coin based Byzantine Randomized Consensus algorithms [9], in which correct processes choose their next estimate by tossing a coin, if they do not manage to reach agreement in a given round. Here we are relying on the random schedule instead of the random coin. However, as we show later, in our model, consensus is solvable for settings where it is impossible for standard asynchrony even when equipped with a common coin.
@@
Crash-fault tolerant consensus. A similar approach can be used to solve crash-fault tolerant consensus with $n=2f+1$, with the same guarantees of deterministic safety, as well as termination almost surely. Our algorithm uses the same round-based structure in Algorithm 1 and a modified Round procedure, that does not require reliable broadcast, and employs standard point-to-point messages instead. The main intuition is similar to our Byzantine algorithm: the random scheduler eliminates the need for a common coin by ensuring that correct processes eventually deliver messages in a favorable order which leads them to agree and thus terminate. Appendix C gives our algorithm and proves its correctness.
4 $n=2f+1$: Deterministic Termination, Safety with High Probability